OpenStudy (anonymous):

integrate 1/sqrt (1-x^2)

OpenStudy (anonymous):

\[\cos^2\theta+\sin^2\theta=1\]

OpenStudy (anonymous):

put x=sin theta

OpenStudy (anonymous):

it is just complicated to write for me..so i will simplify it

OpenStudy (anonymous):